# Group Representations of Lorentz Transformations Extended to Superluminal Observers

### v2

@@ Abstract
where $\mathbb{Z}_{2}$ is $\{$-1,1}\}, the nontrivial automorphism $\text{Ad}_{\Lambda_{\infty}}$ is conjugation by $\Lambda_{\infty}(\theta,\phi)$ and conjugation with matrices corresponding to different values $(\theta^{\prime},\phi^{\prime})$ produces equivalent group extensions. $\mathcal{L}_{\mathrm{ext}}$ has the same identity component $\mathrm{SO}(3,1)^{+}_{\uparrow}$ as the usual Lorentz group but components $\pm\Lambda_{\infty}$ in place of parity and time–reversal of $\mathrm{O}(3,1)$. Each transformation in this extended group preserves the Minkowski form up to sign and can exchange timelike and spacelike directions. Passing to the extended Poincaré group $\mathcal{P}_{\mathrm{ext}}=\mathbb{R}^{4}\rtimes\mathcal{L}_{\mathrm{ext}}$, we classify unitary irreducible representations (UIRs). The action of $\mathcal{L}_{\mathrm{ext}}$ on momentum space merges the timelike and spacelike $\mathrm{SO}(3,1)^{+}_{\uparrow}$–orbits with invariants $p^{\mu}p_{\mu}=\pm M^{2}$ into a single orbit $|p^{\mu}p_{\mu}|=M^{2}$, while merges light-like $\mathrm{SO}(3,1)^{+}_{\uparrow}$-orbits with $p_{0}<0$ and $p_{0}>0$. The resulting induced UIRs for $M\neq 0$ restrict to the ordinary Poincaré subgroup as a multiplicity–one direct sum of a massive forward ($p_{0}>0,p^{\mu}p_{\mu}>0$), a massive backward ($p_{0}<0,p^{\mu}p_{\mu}>0$) and a tachyonic ($p^{\mu}p_{\mu}<0$) Wigner’s UIR, while for $M=0$ as superpositions of forward and backward massless Wigner’s UIRs depending on a two valued parameter. We then derive wave equations corresponding to solutions of the Casimir eigenvalue problem of Poincarè algebra obtained differentiating the above representations. This set of equations contains all the wave equations known to date in quantum field theory together with new wave equations describing tachyonic behaviour and a new class of massless representations. We finally show that tachyonic wave functions provide a relevant representation theoretic tool for interpretation of parity–violation phenomena in quantum field theory
@@ 3.2 Classification of UIRs and Relationship with Wigner’s Representations
    $\pi_{\text{massless}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}=0\}$ (38)
@@ 3.2.2 UIRs Corresponding to Lightlike Orbits
corresponding to forward and backward light–cones ($p_{0}^{0}>0$ and $p_{0}^{0}<0$). They are both induced from unitary irreps of the Euclidean group $ISO(2)$ (helicity or continuous–spin class) and are unitarily equivalent as representations of $\mathcal{P}_{0}$.
In the extended Lorentz group $\mathcal{L}_{\mathrm{ext}}$ the matrix $\Lambda_{-\infty}$ (see Section 2) satisfies $\Lambda_{-\infty}p_{0}=p_{0},$ hence it lies in the stability subgroup of the chosen representative. Consequently, the geometric little group at $p_{0}$ in $\mathcal{L}_{\mathrm{ext}}$ is
@@
Let $k_{0}(p)\in SO(3,1)^{+}_{\uparrow}$ be a section such that $k_{0}(p)\,p_{0}=p$, and define the corresponding Wigner element
$s_{0}(h,p)\;=\;k_{0}(p)^{-1}\,h\,k_{0}(h^{-1}p),\qquad h\in SO(3,1)^{+}_{\uparrow}.$ (66)
This lies in $ISO(2)$ for each $p\in\mathcal{O}_{0}$, and in the extended group it is normalized by $\Lambda_{-\infty}$, so that $\Lambda_{-\infty}s_{0}(h,p)\Lambda_{-\infty}^{-1}\in ISO(2).$
Let $U_{0}$ be a unitary representation of the little group $ISO(2)\rtimes\{I,\Lambda_{-\infty}\}$ associated with either helicity or continuous–spin. The induced Poincaré action on wavefunctions $\psi:\mathcal{O}_{0}\to\mathcal{H}_{0}$ has the usual form
@@
thus interchanging the forward and backward light–cones. In particular $\Lambda_{\infty}$ does not belong to the geometric stabilizer of $p_{0}$. However, at the level of representations of $\mathcal{P}_{0}$ the massless UIRs $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ are equivalent, hence there exists a unitary intertwiner $C$ on the Wigner space $\mathcal{H}_{0}$ such that
$C\;\pi^{\text{fwd}}(n)\;C^{-1}\;=\;\pi^{\text{fwd}}\bigl(\Lambda_{\infty}^{-1}n\Lambda_{\infty}\bigr)\;=\;\pi^{\text{bwd}}(n),\qquad\forall n\in\mathcal{P}_{0}.$ (69)
@@
$\bar{U}(\Lambda_{\infty})\Psi\,](p):=\bigl(C\,\psi_{\text{bwd}}(\Lambda_{\infty}^{-1}p),\;C^{-1}\,\psi_{\text{fwd}}(\Lambda_{\infty}^{-1}p)\bigr),$ (73)
@@
The only further constraint on $\bar{U}(\Lambda_{\infty})$ is that $\bar{U}(\Lambda_{\infty})^{2}=\mathbb{I}.$ This restricts $C$ in (73) to a unitary with $C^{2}=\mathbb{I}$ on $\mathcal{H}_{0}$, so that its eigenvalues are $\pm 1$. Different choices of the sign pattern for the eigenvalues of $C$ lead to two inequivalent massless UIRs of the extended group for each standard Wigner class (helicity or continuous–spin). In this sense, the lightlike orbit supports a doublet representation
@@
where $\varepsilon$ labels the choice of sign in the representation of the discrete $\mathbb{Z}_{2}$ coming from the Mackey stabilizer of the massless UIR under the action of $Z_{(\theta,\phi)}$.

### v3

@@ Abstract
where $\mathbb{Z}_{2}$ is $\{$-1,1}\}, the nontrivial automorphism $\text{Ad}_{\Lambda_{\infty}}$ is conjugation by $\Lambda_{\infty}(\theta,\phi)$ and conjugation with matrices corresponding to different values $(\theta^{\prime},\phi^{\prime})$ produces equivalent group extensions. $\mathcal{L}_{\mathrm{ext}}$ has the same identity component $\mathrm{SO}(3,1)^{+}_{\uparrow}$ as the usual Lorentz group but components $\pm\Lambda_{\infty}$ in place of parity and time–reversal of $\mathrm{O}(3,1)$. Each transformation in this extended group preserves the Minkowski form up to sign and can exchange timelike and spacelike directions. Passing to the extended Poincaré group $\mathcal{P}_{\mathrm{ext}}=\mathbb{R}^{4}\rtimes\mathcal{L}_{\mathrm{ext}}$, we classify unitary irreducible representations (UIRs). The action of $\mathcal{L}_{\mathrm{ext}}$ on momentum space merges the timelike and spacelike $\mathrm{SO}(3,1)^{+}_{\uparrow}$–orbits with invariants $p^{\mu}p_{\mu}=\pm M^{2}$ into a single orbit $|p^{\mu}p_{\mu}|=M^{2}$, while merges light-like $\mathrm{SO}(3,1)^{+}_{\uparrow}$-orbits with $p_{0}<0$ and $p_{0}>0$. The resulting induced UIRs for $M\neq 0$ restrict to the ordinary Poincaré subgroup as a multiplicity–one direct sum of a massive forward ($p_{0}>0,p^{\mu}p_{\mu}>0$), a massive backward ($p_{0}<0,p^{\mu}p_{\mu}>0$) and a tachyonic ($p^{\mu}p_{\mu}<0$) Wigner’s UIR, while for $M=0$ as direct sum of forward and backward massless Wigner’s UIRs depending on a two valued parameter. We then derive wave equations corresponding to solutions of the Casimir eigenvalue problem of Poincarè algebra obtained differentiating the above representations. This set of equations contains all the wave equations known to date in quantum field theory together with new wave equations describing tachyonic behaviour and a new class of massless representations. We finally show that tachyonic wave functions provide a relevant representation theoretic tool for interpretation of parity–violation phenomena in quantum field theory
@@ 3.2 Classification of UIRs and Relationship with Wigner’s Representations
    $\pi_{\text{massless}}^{\text{fwd}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}=0\;\;p^{0}>0\}$ (38)
    $\pi_{\text{massless}}^{\text{bwd}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}=0\;\;p^{0}<0\}$ (39)
@@ 3.2.2 UIRs Corresponding to Lightlike Orbits
corresponding to forward and backward light–cones ($p_{0}^{0}>0$ and $p_{0}^{0}<0$). They are both induced from unitary irreps of the Euclidean group $ISO(2)$ (helicity or continuous–spin class) and are inequivalent as representations of $\mathcal{P}_{0}$.
In the extended Lorentz group $\mathcal{L}_{\mathrm{ext}}$ the matrix $\Lambda_{-\infty}=-\Lambda_{\infty}$ satisfies $\Lambda_{-\infty}p_{0}=p_{0},$ hence it lies in the stability subgroup of the chosen representative. Consequently, the geometric little group at $p_{0}$ in $\mathcal{L}_{\mathrm{ext}}$ is
@@
Let $U_{0}$ be a unitary representation of the little group $ISO(2)$ associated with either helicity or continuous–spin. The induced Poincaré action on wavefunctions $\psi:\mathcal{O}_{0}\to\mathcal{H}_{0}$ has the usual form
@@
thus interchanging the forward and backward light–cones. In particular $\Lambda_{\infty}$ does not belong to the geometric stabilizer of $p_{0}$. However, at the level of massless UIRs of $\mathcal{P}_{\text{ext}}$, $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ are equivalent since the unitary operator representing $U(\Lambda_{\infty})$ is thus an intertwiner between $\pi_{\text{fwd}}$ and $\pi_{\text{bwd}}$
$U(\Lambda_{\infty})\;\pi^{\text{fwd}}(n)\;U(\Lambda_{\infty})^{-1}\;=\;\pi^{\text{fwd}}\bigl(\Lambda_{\infty}^{-1}n\Lambda_{\infty}\bigr)\;=\;\pi^{\text{bwd}}(n),\qquad\forall n\in\mathcal{P}_{0}.$ (69)
@@
$\bar{U}(\Lambda_{\infty})\Psi\,](p):=\bigl(U(\Lambda_{\infty})\,\psi_{\text{bwd}}(\Lambda_{\infty}^{-1}p),\;U(\Lambda_{\infty})^{-1}\,\psi_{\text{fwd}}(\Lambda_{\infty}^{-1}p)\bigr),$ (73)
@@
The only further constraint on $\bar{U}(\Lambda_{\infty})$ is that $\bar{U}(\Lambda_{\infty})^{2}=I.$ This restricts $U(\Lambda_{\infty})$ in (73) to a unitary with $U(\Lambda_{\infty})^{2}=\mathbb{I}$ on $\mathcal{H}_{0}$, so that its eigenvalues are $\pm 1$. Different signs for the eigenvalues of $U(\Lambda_{\infty})$ lead to two inequivalent massless UIRs of the extended group for each standard Wigner class (helicity or continuous–spin). In this sense, the lightlike orbit supports a doublet representation
@@
where $\varepsilon$ labels the choice of sign in the representation of the extended little group of lightlike momenta representatives.